December 14, 2016 - 4-Laser/LLRF Installation Meeting

From Ciswikidb
Jump to: navigation, search

We will meet Wednesday Dec 14 @ 10am in TL-1227 for final updates.

Meeting Minutes

Schedule
Dec. 21 : benchmark 4 laser w/ present LLRF, laser amplifiers software upgrade for current control, turn amps off and safe laser for RF upgrade
Jan 3-13 : install new RF control hardware and cabling, install software and update screens, connect to lasers, turn lasers back ON, complete INJ SEG HCO
Jan 17-20 : system commissioning/float, request INJ SEG=BP for beam checkout, add fixes or final improvements, complete 4-laser upgrade, complete INJ/NL HCO
Jan 23-25 : request INJ/NL=BP full energy injector setup, tests w/ 4 beams @ injector
Jan 26 : ready for beam operations
High Level Notables
No spares of critical components, budget for spares being estimated
Existing 3-LLRF will be retained as roll-back, rack chassis left if possible or stored, VME cards retained as-is
Tomasz will submit LLRF ATLIS for comments and tracking of activity
Hardware installation begins first thing Jan. 3; in good shape, but still awaiting amplifiers due before holiday break
Firmware progressing well w/ amplitude and phase set point by EPICS, resolution and stability good (detailed report below)
Additional EPICS integration in January, will be tight to complete (Scott/Rama to assess by holiday, raise flag if problematic); will address how to assign 499/250 to single phase shifter
All on same page that no phase jumps anticipated in running system, however, until MO upgrade too lower freq. expect phase jump w/ power outage, cable change, need procedure for checkout/Ops

Status Updates

Joe - Schedule
Beam Study 12/21 8am-Noon will benchmark e- bunch length and laser power and record archived high level RF parameters (power/phase)
media:161214_LLRF_INSTALL.pptx
Winter SAD
Tomasz - LLRF upgrade status
Rama/Chad - Firmware
progress on firmware
1.PC104 breakout board was tested and is being used with Scott's software and the communication between EPICS and FPGA looks good
2.FPGA is talking to DAC (4-channel), ADC(4 channel) and PC104.
3.Tone mode functionality needed for January is now working. Magnitude and Phase can be set to 4 DAC channels individually from EPICS PVs
4.Magnitude and Phase measured from ADC channels is displayed in EPICS
5.ADC and DAC have enough resolution for control when needed (0.2 degree resolution for phase and 0.5% for magnitude), so there is no need for an extra clock board
tasks in progress for January
1.replacing 4 channel ADC board with 8 channel ADC board. Not needed for January, but will install once testing is done in next couple of days. This will give us the ability to scan all the channels (4 lasers, 70 MHz, 35 MHz)
2.work on screens for OPS
3.procedure for correcting phase, if the phase changes in either 35 MHz or 70 MHz (EPICS level)
4.testing with RF board once the chassis is assembled (in progress)
5.continuing firmware development
Scott - Software
Dave/Manuel - Hardware
Dave - Documentation/Drawings
Michael/John - Current control and remote operation of laser amplifiers
Michele - 4-laser HLAPP
Application ready to test